An expert shot hits the target 90% of the time. what is the probability that he will miss the target for the first time on the 10th shot?
Answer by Theo(3504)

The probability that he will hit the target on each shot is 9/10 = 90%.

This means that the probability that he will miss the target on each shot is 1/10 = 10%.

The probability that he gets 9 his in a row and 1 miss on the 10th shot would be:

(9/10) * (9/10) * (9/10) * (9/10) * (9/10) * (9/10) * (9/10) * (9/10) * (9/10) * (1/10) which is equivalent to:

(9/10)^9*(1/10)^1 = .038742049 = 3.87% rounded to the nearest hundredth of a percent.